Please show work

1) A trader entered into a long S&P index futures contract when the futures price is 3,112 and closed it out when the futures price is 3,075. What is his gain or loss in dollars? The contract multiplier is $250. Write your answer in the format as 10000 or -10000, no dollar sign, no comma.

2) A trader entered into 2 short British Pound futures contracts @ $1.6 per GBP, and closed out his position @ $1.71 per GBP. What is his gain or loss in USD? The size of the contract is 62,500 GBP. Write your answer in the format of 10000 or -10000, no comma, no dollar sign.

Please show work
Answer #1

Solution:-

Contract multiplier is essentially the profit or loss on each point gained or lost in the futures contract.

(1) No. of points gained / (lost) on futures contract = 3,075-3,112 = -37 points

Loss on trade = points lost*contract multiplier = -37*250 = -9250

(2) The trader shorted 2 contracts.

Selling price per contract ($)= GBP62,500*1.6 = $100,000

Purchase price per contract ($)= GBP62,500*1.71 = $106,875

$ profit / (loss) on two contracts = (100,000-106,875)*2 = -13750